Solve (Factoring)
3(x − 1) = 9 −x

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

3x-3=9-x
4x=12
x=12/4
x=3

Answer by stanbon(75887)

3(x − 1) = 9 −x
----
No factoring is involved.
-----
3x-3 = 9-x
4x = 12
x = 3
============
Cheers,
Stan H.
======================